# A generalization of Zalcman’s Lemma on Complex Lie Groups

### v1

@@ Abstract.
Normal families; Zalcman’s lemma; Marty’s criteria; Heuristic principle; Complex Lie groups
@@ 1. Introduction
Zalcman’s work was inspired by the result of Lohwater and Pommerenke [10], which establishes a well-known heuristic principle showing that Bloch’s principle [2] is valid to a certain Picard type property. Due to the importance of Zalcman’s lemma, many authors have been working on its generalizations. A beautiful work on this lemma is due to Pang [13, 14] who proved a stronger version with a real parameter, which thus shows that Bloch’s principle holds for some Picard type properties related to derivatives. More details for this lemma and its applications in one complex variable refer to [4, 9, 16, 20, 23] and their references.
@@
For $N=\mathbb{P}^{1}(\mathbb{C}),$ Theorem B was generalized by Chang, Xu and Yang [5], Dovbush [7], Yang and Pang [21], etc. More details see [6] and its references.
@@
Afterwards, we shall state a Robinson-Zalcman type heuristic principle of holomorphic families of a Picard type property on a complex Lie group. Let $g$ be a holomorphic mapping on a domain $D$ in a complex manifold. Denote by $P$ a property. Notation $(g,D)\in P$ denotes that $g$ is of $P$ on $D,$ notation $(g,x)\in\neg P$ denotes that $g$ is not of $P$ in any neighborhood of $x$ in $D,$ and notation $(g,D)\in\neg P$ denotes that $(g,x)\in\neg P$ for all $x\in D.$
@@ 3. Proofs and Picard Theorem
Let $\mathscr{C}(M,N)$ denote the set of all continuous mappings between complex manifolds $M,N.$ Equip $\mathscr{C}(M,N)$ with the compact-open topology. A family $\mathscr{F}\subset\mathscr{C}(M,N)$ is called normal, if and only if each sequence of $\mathscr{F}$ contains a subsequence which is either relatively compact in $\mathscr{C}(M,N)$ or compactly divergent (Definition 1.1, [19]). Let $\mathscr{H}(M,N)\subset\mathscr{C}(M,N)$ denote the set of all holomorphic mappings between $M,N.$ Anyway, once a distance function is chosen on $N,$ a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{C}(M,N)$ converges to an $f\in\mathscr{C}(M,N)$ if and only if $\{f_{j}\}_{j=1}^{+\infty}$ converges uniformly on compact subsets to $f.$ Hence, a basic fact derived from Weierstrass theorem shows that $\mathscr{H}(M,N)$ is closed in $\mathscr{C}(M,N).$ When $N$ is a compact Hermitian manifold, by a result of Wu (Lemma 1.1, [19]), $\mathscr{F}$ is normal on $M$ whenever $\mathscr{F}$ is equicontinuous on $M.$
To prove Theorem 1.1, we extend Marty’s criteria [12] for normality. Let $f:M\to N$ be a holomorphic mapping between Hermitian manifolds $M,N.$ The operator norm $\|df\|$ of differential $df$ is defined by
@@
We provide a slight extension of Marty’s criteria as follows.

### v2

@@ Abstract.
Normal families; Zalcman’s lemma; Marty’s criterion; Heuristic principle; Complex Lie groups
@@ 1. Introduction
Zalcman’s work was inspired by the result of Lohwater and Pommerenke [10], which establishes a well-known heuristic principle showing that Bloch’s principle [2] is valid to a certain Picard type property. Due to the importance of Zalcman’s lemma, many authors have been working on its generalizations. A beautiful work on this lemma is due to Pang [13, 14] who proved a stronger version with a real parameter, which thus shows that Bloch’s principle holds for some Picard type properties related to derivatives. More details for this lemma and its applications in one complex variable refer to [4, 9, 16, 20, 23] and references therein.
@@
For $N=\mathbb{P}^{1}(\mathbb{C}),$ Theorem B was generalized by Chang, Xu and Yang [5], Dovbush [7], Yang and Pang [21], etc. More details refer to [6] and references therein.
@@
Afterwards, we shall state a Robinson-Zalcman type heuristic principle of holomorphic families of a Picard type property on a complex Lie group. Let $g$ be a holomorphic mapping on a domain $D$ in a complex manifold. Denote by $P$ a property. Write $(g,D)\in P,$ if $g$ is of $P$ on $D;$ $(g,x)\in\neg P,$ if $g$ is not of $P$ in any neighborhood of $x$ in $D;$ and $(g,D)\in\neg P,$ if $(g,x)\in\neg P$ for all $x\in D.$
@@ 3. Proofs and Picard Theorem
Let $\mathscr{C}(M,N)$ denote the set of all continuous mappings between complex manifolds $M,N.$ Equip $\mathscr{C}(M,N)$ with the compact-open topology. A family $\mathscr{F}\subset\mathscr{C}(M,N)$ is called normal, if and only if each sequence of $\mathscr{F}$ contains a subsequence which is either relatively compact in $\mathscr{C}(M,N)$ or compactly divergent (Definition 1.1, [19]). Let $\mathscr{H}(M,N)\subset\mathscr{C}(M,N)$ denote the set of all holomorphic mappings between $M,N.$ Anyway, once a distance function is chosen on $N,$ a sequence $\{f_{j}\}_{j=1}^{+\infty}\subset\mathscr{C}(M,N)$ converges to an $f\in\mathscr{C}(M,N)$ if and only if $\{f_{j}\}_{j=1}^{+\infty}$ converges uniformly on compact subsets to $f.$ Hence, a basic fact derived from Weierstrass theorem shows that $\mathscr{H}(M,N)$ is closed in $\mathscr{C}(M,N).$ When $N$ is a compact Hermitian manifold, using Ascoli-Arzelà theorem or a result of Wu (Lemma 1.1, [19]), $\mathscr{F}$ is normal on $M$ as long as $\mathscr{F}$ is equicontinuous on $M.$
To prove Theorem 1.1, we extend Marty’s criterion [12] for normality. Let $f:M\to N$ be a holomorphic mapping between Hermitian manifolds $M,N.$ The operator norm $\|df\|$ of differential $df$ is defined by
@@
We provide a slight extension of Marty’s criterion as follows.